# Asymptotic theory of in-context learning by linear attention

### v1

@@ Appendix B The Stieltjes Transforms of Wishart Ensembles
$\tau\chi^{\prime}_{0}=\frac{1}{d}\operatorname{tr}\left[(E_{\mathrm{tr}}+\xi_{0}I)^{-1}\Big{(}\chi^{\prime}_{0}E_{\mathrm{tr}}-\frac{(1+\chi_{0})^{2}}{\tau}B_{\mathrm{test}}\Big{)}(E_{\mathrm{tr}}+\xi_{0}I)^{-1}E_{\mathrm{tr}}\right].$ (B.17)
@@
$\frac{\tau\chi^{\prime}_{0}}{(1+\chi_{0})^{2}}$ $\simeq\frac{\frac{1}{d}\operatorname{tr}\left(B_{\mathrm{test}}[(E_{\mathrm{tr}}+\xi I)^{-1}-\xi(E_{\mathrm{tr}}+\xi I)^{-2}]\right)}{1-2\xi\operatorname{tr}(E_{\mathrm{tr}}+\xi I)^{-1}/d+\xi^{2}\operatorname{tr}(E_{\mathrm{tr}}+\xi I)^{-2}/d-\tau},$ (B.18)
where we have used (B.16) to replace $\xi_{0}$ in (B.17) by $\xi$, with the latter being the solution to the self-consistent equation in (SI.5.6). Using the decomposition in (B.7) and following similar arguments that allowed us to simplify (B.10) to (B.13), we can check that
$\frac{1}{d}\operatorname{tr}(E_{\mathrm{tr}}+\xi I)^{-1}\simeq\frac{1}{d}\operatorname{tr}S(E_{\mathrm{tr}}+\xi I)^{-1}S^{\top}\simeq\frac{1}{d}\operatorname{tr}F\left(\frac{1+\rho}{\alpha}+\xi\right)\simeq\mathcal{M}_{\kappa}\left(\frac{1+\rho}{\alpha}+\xi\right),$ (B.19)
@@
$\frac{1}{d}\operatorname{tr}(E_{\mathrm{tr}}+\xi I)^{-2}\simeq\frac{1}{d}\operatorname{tr}S(E_{\mathrm{tr}}+\xi I)^{-2}S^{\top}\simeq\frac{1}{d}\operatorname{tr}F^{2}\left(\frac{1+\rho}{\alpha}+\xi\right)\simeq-\mathcal{M}^{\prime}_{\kappa}\left(\frac{1+\rho}{\alpha}+\xi\right),$ (B.20)
where $S$ is a $d\times(d+1)$ matrix obtained by removing the last row of $I_{d+1}$, and $\mathcal{M}_{\kappa}(\cdot)$ is the function defined in (B.3). Substituting (B.19) and (B.20) into (B.18) yields
$\frac{\tau\chi^{\prime}_{0}}{(1+\chi_{0})^{2}}$ $\simeq\frac{\frac{1}{d}\operatorname{tr}\left(B_{\mathrm{test}}[(E_{\mathrm{tr}}+\xi I)^{-1}-\xi(E_{\mathrm{tr}}+\xi I)^{-2}]\right)}{1-2\xi\mathcal{M}_{\kappa}\left(\frac{1+\rho}{\alpha}+\xi\right)-\xi^{2}\mathcal{M}^{\prime}_{\kappa}\left(\frac{1+\rho}{\alpha}+\xi\right)-\tau}.$ (B.21)

### v2

@@ Acknowledgments
$\tau\chi^{\prime}_{0}=\frac{1}{d}\operatorname{tr}\left[(E_{\mathrm{tr}}+\xi_{0}I)^{-1}\Big{(}\chi^{\prime}_{0}E_{\mathrm{tr}}-\frac{(1+\chi_{0})^{2}}{\tau}B_{\mathrm{test}}\Big{)}(E_{\mathrm{tr}}+\xi_{0}I)^{-1}E_{\mathrm{tr}}\right].$ (SI.8.17)
@@
$\frac{\tau\chi^{\prime}_{0}}{(1+\chi_{0})^{2}}$ $\simeq\frac{\frac{1}{d}\operatorname{tr}\left(B_{\mathrm{test}}[(E_{\mathrm{tr}}+\xi I)^{-1}-\xi(E_{\mathrm{tr}}+\xi I)^{-2}]\right)}{1-2\xi\operatorname{tr}(E_{\mathrm{tr}}+\xi I)^{-1}/d+\xi^{2}\operatorname{tr}(E_{\mathrm{tr}}+\xi I)^{-2}/d-\tau},$ (SI.8.18)
where we have used (SI.8.16) to replace $\xi_{0}$ in (SI.8.17) by $\xi$, with the latter being the solution to the self-consistent equation in (SI.5.6). Using the decomposition in (SI.8.7) and following similar arguments that allowed us to simplify (SI.8.10) to (SI.8.13), we can check that
$\frac{1}{d}\operatorname{tr}(E_{\mathrm{tr}}+\xi I)^{-1}\simeq\frac{1}{d}\operatorname{tr}S(E_{\mathrm{tr}}+\xi I)^{-1}S^{\top}\simeq\frac{1}{d}\operatorname{tr}F\left(\frac{1+\rho}{\alpha}+\xi\right)\simeq\mathcal{M}_{\kappa}\left(\frac{1+\rho}{\alpha}+\xi\right),$ (SI.8.19)
@@
$\frac{1}{d}\operatorname{tr}(E_{\mathrm{tr}}+\xi I)^{-2}\simeq\frac{1}{d}\operatorname{tr}S(E_{\mathrm{tr}}+\xi I)^{-2}S^{\top}\simeq\frac{1}{d}\operatorname{tr}F^{2}\left(\frac{1+\rho}{\alpha}+\xi\right)\simeq-\mathcal{M}^{\prime}_{\kappa}\left(\frac{1+\rho}{\alpha}+\xi\right),$ (SI.8.20)
where $S$ is a $d\times(d+1)$ matrix obtained by removing the last row of $I_{d+1}$, and $\mathcal{M}_{\kappa}(\cdot)$ is the function defined in (SI.8.3). Substituting (SI.8.19) and (SI.8.20) into (SI.8.18) yields
$\frac{\tau\chi^{\prime}_{0}}{(1+\chi_{0})^{2}}$ $\simeq\frac{\frac{1}{d}\operatorname{tr}\left(B_{\mathrm{test}}[(E_{\mathrm{tr}}+\xi I)^{-1}-\xi(E_{\mathrm{tr}}+\xi I)^{-2}]\right)}{1-2\xi\mathcal{M}_{\kappa}\left(\frac{1+\rho}{\alpha}+\xi\right)-\xi^{2}\mathcal{M}^{\prime}_{\kappa}\left(\frac{1+\rho}{\alpha}+\xi\right)-\tau}.$ (SI.8.21)